# Dockformer: A transformer-based molecular docking paradigm for large-scale virtual screening

### v2

@@ V CONCLUSION
With the aid of artificial intelligence technologies, large-scale compound libraries can be explored to pursue promising drug candidates with higher diversity and stronger binding strengths, which may improve the hit rates to target specific proteins of interest. Following this perspective, an in silico molecular docking algorithm named Dockformer, whose end-to-end framework enables improvements in docking accuracy and screen efficiency simultaneously, was proposed in this study. However, although existing computational methods have undergone tremendous advances, detecting the biological interactions between proteins and ligands remains challenging because of the scarcity of available training data, physical implausibility, and false-positives of hit identifications. In addition, the Chemical Universe Databases will contain trillions of compounds in the coming years, requiring more efficient high-performance screening methods to search such extremely large regions of the chemical space. De novo drug design methods may be an alternative to docking algorithms to skip the computationally expensive screening process on the basis of generative models and deep reinforcement learning approaches.

### v4

@@ V CONCLUSION
This study proposes Dockformer for LSVS, which integrates multimodal fusion, positional encoding and end-to-end architecture. Compared with conventional DL-based algorithms, these advanced components enable Dockformer to improve docking accuracy and accelerate the docking process significantly. In addition, Dockformer showcases its potential to expedite drug discovery efforts by efficiently screening large molecular libraries and precisely identifying compounds with similar interaction patterns to known inhibitors. As a robust and reliable protein-ligand docking approach, Dockformer holds promise for significantly reducing the development cycle and cost of drug design.
With the aid of artificial intelligence technologies, large-scale compound libraries can be explored to pursue promising drug candidates with higher diversity and stronger binding strengths, which may improve the hit rates to target specific proteins of interest. Following this perspective, an in silico molecular docking algorithm named Dockformer, whose end-to-end framework enables improvements in docking accuracy and screen efficiency simultaneously, is proposed in this study. However, although existing computational methods have undergone tremendous advances, detecting the biological interactions between proteins and ligands remains challenging because of the scarcity of available training data, physical implausibility, and false-positives of hit identifications. In addition, the Chemical Universe Databases will contain trillions of compounds in the coming years, requiring more efficient high-performance screening methods to search such extremely large regions of the chemical space. De novo drug design methods may be an alternative to docking algorithms to skip the computationally expensive screening process on the basis of generative models and deep reinforcement learning approaches. The source code of Dockformer is available at https://zenodo.org/records/12792385.